# How Internal Structure Shapes the Metallicity of Giant Exoplanets

### v1

@@ 3.1 Structural properties
Figure 2 shows the retrieved curves for CE, FM, and three DC structures that assume a different extent of the dilute core ($m_{\rm dilute}=$ 0.25, 0.5 or 0.75). Although the analysis is performed for the full sample, we first focus on three planets with representative masses: K2-287b (0.315 $M_{\rm J}$), TOI-1811b (0.972 $M_{\rm J}$), and TOI-5153b (3.26 $M_{\rm J}$). In all cases, CE yields a lower bulk metallicity than FM. We verified this offset by reproducing the approach of Howard et al. (2025b); the results are consistent with our pipeline (see Appendix B). The physical driver is that, in FM, heavy elements mixed into the envelope and atmosphere increase the mean molecular weight and, critically, the Rosseland mean opacity in the radiative zone. The higher opacity slows cooling, and keeps radii larger at fixed age; matching the observed radius then requires a larger heavy-element mass than in the CE case. Howard et al. (2025b) likewise found FM evolution tracks to be systematically warmer, especially at early ages.The opacity effect is the main contributor in FM structures for higher expected bulk metallicities.
@@
When examining the DC curves at different $m_{\rm dilute}$ values (Fig. 2), we find that DC models generally underestimate $Z$ compared to CE and FM, and their solutions more closely resemble CE than FM. No systematic distinction or trend is evident across the three different $m_{\rm dilute}$ values tested, as the model-to-model differences (and their intrinsic uncertainties) remain smaller than the observational errors. Since these DC structures exclude atmospheric metallicity ($Z_{\rm atm}=0$), the observed effects are driven solely by the internal distribution of heavy elements. Introducing a dilution gradient, rather than a sharp core boundary, redistributes heavy material into outer layers, increasing local density and pressure. This accelerates planetary contraction over time, meaning that a smaller heavy-element mass is required to reproduce the observed radius and age. Nonetheless, such an effect appears to be of lower magnitude compared to the FM-CE estimates difference.
@@
After testing different $m_{\rm dilute}$ structures, we now examine dilute core models with atmospheric metallicities set to one and three times the stellar value (DCA and DCA3, respectively). Results for the three representative planets are shown in Fig. 3, while the full dataset, including mean values and standard deviations for all structures, is reported in Appendix LABEL:app:sample. Since $Z_{\rm atm}$ (and thus the opacity effect) is the varying parameter, DCA and DCA3 directly probe its effect on $Z$ estimates. As Fig. 3 illustrates, increasing atmospheric metallicity systematically shifts the inferred $Z$ upward. As a result, in the majority of cases DCA expects higher heavy elements in planetary interiors, followed by DCA3 and FM. In the three cases shown, both DCA and DCA3 yield higher $Z$ than CE. For some planets (e.g., HAT-P-12b, TOI-2010b, TOI-2180b), however, DCA results converge with DC models, suggesting that opacity effects are insufficient to alter the estimates. Overall, Fig. 3 confirms that opacity is the primary driver of the retrieved heavy-element abundance, whereas density redistribution alone produces comparatively minor shifts. The CEA (bright-red) curves closely track the DCA estimates (especially for TOI-1811b and TOI-5153b in the figure), indicating that both models recover similar $Z$. Because they use identical opacities and differ only in interior structure (distinct core in CEA vs. composition gradient in DCA), this agreement reinforces that opacity is the dominant lever over density redistribution.
In the absence of additional heat sources, opacity controls the degree of radius inflation in our models. Consequently, the magnitude of the opacity effect is planet-specific and cannot be captured by a single global scaling (reflecting planetary variations between transport, boundary conditions, and composition degeneracies). Furthermore, fixed atmospheric constraints such as a three times $Z_{\star}$ is ruling out possible structural scenarios for low-$Z$, high-$M$ planetary systems, where there are cases where DCA3 and FM models give similar results (like in the case of TOI-5153b in Fig.3), or even in some cases the DCA3 estimates exceed the FM ones (like TOI-4127b or WASP-162b cases that can be fount in the Appendix LABEL:app:sample). An atmospheric enrichment of $3\,Z_{\star}$ therefore creates a bias by the model in pushing metallicities of low estimated $Z$ up. Tightening atmospheric constraints (e.g., on metallicity and thermal structure) therefore becomes crucial for breaking high-$M$ degeneracies.
@@ 3.2 Mass-Metallicity Relation
After analysing the overall behaviour of metallicity estimates across different structural models, we next examined the full dataset of 44 planets for possible population-level trends in the inferred composition. In particular, we investigated the heavy-element mass ($M_{Z}$) and the bulk metallicity ($Z$), and how they scale with planetary mass ($M$).
@@
Figure 4: Relationship between the retrieved heavy element mass $M_{Z}$ and planetary mass $M$, shown for the four main structures tested (CE, FM, DCA, and DCA3). Shown as dashed lines in gray gradient are the different curves showing how the trend looks for fixed planetary compositions (0%, 50%, 90%, and 99% gas composition respectively). Bayesian fits have been shown using a power law function $M_{Z}=\beta\,M^{\alpha}$. Median values of $\beta$ are respectively 59.57 for FM, 56.02 for DCA3, 43.11 for DCA, and 35.9 for CE. Values for $\alpha$ are 0.653 for FM, 0.718 for DCA3, 0.700 for DCA, and 0.667 for CE. Posterior uncertainties $\sigma_{\alpha}$ and $\sigma_{\beta}$ are respectively 0.068 and 4.5 for FM, 0.074 and 4.3 for DCA3, 0.098 and 4.4 for DCA, and 0.113 and 4.5 for CE.
@@
To assess these relations, we performed a Bayesian regression. The parameters of the power-law $M_{Z}\,[M_{\oplus}]=\beta\,M\,[M_{J}]^{\alpha}$ are estimated using a Markov chain Monte Carlo (MCMC, Python package emcee (Foreman-Mackey et al., 2013), following Müller & Helled (2023) and Howard et al. (2025b)). Posterior distributions of the slope ($\alpha$) and intercept ($\beta$) were sampled from 5,000 chains, taking into account $1\sigma$ uncertainties in both axes. Medians and standard deviations have been derived from these posteriors are shown in Fig. 4 for the four structural models (FM, DCA3, DCA, CE). All models reveal a positive sloping between $M_{Z}$ and $M$, with similar slopes but offsets in intercept depending on the assumed structure. FM systematically predicts the highest $M_{Z}$, while CE predicts the lowest (with $\beta\approx 36$). This scaling is broadly consistent with the relation reported by Thorngren et al. (2016), Howard et al. (2025b), and Chachan et al. (2025). Our results suggest a steeper dependence, with DCA3 yielding $\alpha>0.7\pm 0.07$, only marginally inconsistent with the Thorngren’s relation of $M_{Z}\propto\sqrt{M}$ when uncertainties are considered.
@@
For this reason, an inspection of the reliability of such curves has been done by testing the correlation between $\alpha$ and $\beta$. The covariance analysis shows a weak but positive coupling ($\rho\sim 0.2$ for all the models), indicating that slope and intercept are not fully independent, and that small shifts in one parameter do affect the other. We then also evaluated the fits with a weighted least-squares (WLS) $\chi^{2}$ test. In all cases, $\chi^{2}\gg 1$, meaning the power-law relations do not formally explain the data within the quoted errors, and reliance of such curves should be treated with caution (see Section D for more information about the WSL goodness-of-fit approach.
@@
Figure 5: Relationship between the retrieved heavy element mass $Z$ and planetary mass $M$, shown for the four main structures tested (CE, FM, DCA, and DCA3). Bayesian fits have been shown using a power law function $Z=\beta\,M^{\alpha}$. Median values of $\beta$ are respectively 0.187 for FM, 0.176 for DCA3, 0.136 for DCA, and 0.113 for CE. Values for $\alpha$ are -0.347 for FM, -0.282 for DCA3, -0.3 for DCA, and -0.333 for CE. Posterior uncertainty $\sigma_{\alpha}$ and $\sigma_{\beta}$ are respectively 0.069 and 0.014 for FM, 0.073 and 0.014 for DCA3, 0.097 and 0.015 for DCA, and 0.115 and 0.014 for CE.
@@
Since the weighted least-squares test revealed the limited reliability of the retrieved power-law fits, it is essential to verify whether the apparent slopes of these relations genuinely reflect an underlying correlation within the data. We further tested the correlations of the dataset for each structural assumption using the non-parametric Kendall’s $\tau$ rank test. To account for observational uncertainties, instead of computing the correlation on the data’s mean values, we computed a bootstrap resampling approach was employed: 5,000 synthetic datasets were generated by drawing from the reported error distributions, and $\tau$ was calculated for each realization. The median $\tau$ values and associated p-values are reported in Table 1. For the $_{Z}MM$ relation, nearly all runs returned p-values well below the adopted significance threshold of $p=0.05$, confirming the robustness of the positive correlation.
(a) $M_{z}$ vs. $M$ (b) $Z$ vs. $M$
Model $\boldsymbol{\tau}$ p-value $\boldsymbol{\tau}$ p-value
CE 0.514 $1.20\times 10^{-6}$ 0.273-0.273 $1.00\times 10^{-2}$
DCA 0.612 $7.59\times 10^{-9}$ 0.245-0.245 $2.07\times 10^{-2}$
DCA3 0.660 $4.44\times 10^{-10}$ 0.314-0.314 $3.05\times 10^{-3}$
FM 0.660 $4.44\times 10^{-10}$ 0.397-0.397 $1.79\times 10^{-4}$
@@
Table 1: Kendall’s $\tau$ coefficients and associated p-values for four structural models (CE, DCA, DCA3, FM) across the two tested relationships.
@@
We then examined the relation between $Z$ and $M$. In contrast to the $_{Z}MM$ trend, the ZZMM relation shows a clear negative correlation. The ranking of models remains similar: FM predicts the highest average $Z$, followed by DCA3, DCA, and CE. Compared to the $_{Z}MM$ relation, the data are more scattered, particularly at lower planetary masses. Kendall’s $\tau$ again indicates statistically significant negative correlations.
(Table 1), although the correlation strength is weaker. A portion of the reduced significance likely reflects mathematical coupling from dividing by the xx–axis variable ($Z=M_{Z}/M$), which weakens the apparent correlation and inflates p-values. Median p-values are still generally below 0.05, but with a modest fraction of realizations exceeding the threshold ($\sim 2\%$ for FM and DCA3, and $\sim 10$$11\%$ for DCA and CE). Despite this, the correlations remain statistically compelling across all models.
Figure 6: Top: Four panels using the retrieved method explained in Section 2.3. Retrieved metallicity curves for a DCA model with the assumption of adiabaticity are compared with superadiabatic cases where the $R_{\rho}$ scaling factor is either 0.1 or 0.4. This is shown for four different planetary cases. Bottom: Corresponding temperature profiles for the same four planets, using the median metallicity of the DCA model at a fixed age of 4 Gyr. The different heat transport scenarios show a linear increase of heat retention (and thus deviation from adiabaticity) according to the $R_{\rho}$ value defined in Eq. 3. The $x$-axis shows the bulk metallicity $Z$.
@@
Such a relationship has also been tested with the normalisation to the host star ($ZZ_{\star}$ vs $M$). The results show similar trends (both in best fit and correlation tests) to those obtained for ZZ/$M$. Such an investigation would have been essential if we were artificially inputting anti-correlation given from the star’s metallicity (especially since a model assumption is the stellar composition in the planet’s atmospheric structure). Although recent studies (Buchhave et al., 2014; Müller & Helled, 2025) found trends between the planetary heavy element mass and the chemical composition of their host stars, the dataset used in this study does not show any correlation between $Z_{\star}$ and $M$.
@@ 4 Discussion
Figure 7: Panels showing iterative removal of either smallest mass planets (left) or bigger mass ones (right), and how the correlation coefficient p-values changes as a consequence of that. Colour coded are the four different models. The horizontal line shows the p=0.05 acceptance thershold. Planetary mass removed each time is noted on the x-axis.
@@ 4.1 Non-adiabatic structure
In our models with a dilute core (DC, DCA and DCA3), we assumed an adiabatic temperature gradient. However, the presence of a composition gradient implies that the temperature gradient could differ from adiabatic. To assess the effect of a non-adiabatic temperature gradient on the inferred bulk metallicity, we parameterize the deviation from adiabaticity (likely caused by double-diffusive convection (Debras et al., 2021; Leconte & Chabrier, 2012)) as in Howard et al. (2025a), using:
@@
For a smaller selection of planets, we calculated evolution models using $R_{\rho}=0.1$ and 0.4 and compared it to the adiabatic case ($R_{\rho}=0$). The inferred bulk metallicities $Z$ are shown in Fig. 6. We find that $R_{\rho}=0.1$ and $R_{\rho}=0.4$ lead to increased values of $Z$ by up to 7 % and 35 %, respectively. The super-adiabatic temperature gradients obtained with $R_{\rho}=0.1$ and $R_{\rho}=0.4$ yield hotter interiors compared to the adiabatic case (see bottom panel of Fig. 6) and naturally lead to an increased amount of heavy-element to match the planets’ radii.
However, a clear limitation of our models is that convective mixing is not included. Such a process could affect the thermal evolution and the distribution of heavy elements. Knierim & Helled (2025) showed for a Jupiter-mass planet that neglecting mixing could overestimate the radius by 0.08 $R_{\rm J}$. The values of $R_{\rho}$ we chose were arbitrary and it would be necessary to assess whether dilute cores with such a temperature structure can remain stable against convection. Future studies should further assess the effect of non-adiabaticity on the mass–metallicity relation and the thermal evolution of gas giants.

### v2

@@ 3.1 Structural properties
Figure 2 shows the retrieved curves for CE, FM, and three DC structures that assume a different extent of the DC ($m_{\rm dilute}=$ 0.25, 0.5 or 0.75). Since $m_{\rm dilute}$ remained fixed in the three different cases, $Z_{\rm dilute}$ was used as the adjusting parameter. This setup allowed us to explore the impact of changing the extent of the dilution curve. Although the analysis was performed for the full sample, we first focus on three planets with representative masses: K2-287b (0.315 $M_{J}$), TOI-1811b (0.972 $M_{J}$), and TOI-5153b (3.26 $M_{J}$). In all cases, CE yields a lower bulk metallicity than FM. We verified this offset by reproducing the approach of Howard et al. (2025b); the results are consistent with our pipeline (see Appendix B). The physical driver is that, in FM, heavy elements mixed into the envelope and atmosphere increase the mean molecular weight and, critically, the Rosseland mean opacity in the radiative zone. The higher opacity slows cooling and keeps radii larger at fixed age; matching the observed radius then requires higher heavy-element mass than in the CE case. Howard et al. (2025b) also found that FM evolution tracks are systematically warmer, especially at younger ages. The opacity effect is the main contributor in FM structures for higher expected bulk metallicities.
@@
When examining the DC curves for different $m_{\rm dilute}$ values (Fig. 2), we find that DC models generally have lower values of $Z$ compared to CE and FM models, and their solutions more closely resemble CE than FM. No systematic distinction or trend is evident across the three different $m_{\rm dilute}$ values tested, as the model-to-model differences (and their intrinsic uncertainties) remain smaller than the observational errors. Since these DC structures exclude atmospheric metallicity ($Z_{\rm atm}=0$), the observed effects are driven solely by the internal distribution of heavy elements. Introducing a gradient, rather than a sharp core boundary, redistributes heavy elements into outer layers, increasing the local density and pressure. This results in a lower radius over time, because a smaller heavy-element mass is required to reproduce the observed radius and age. Nonetheless, such an effect appears to be of lower magnitude compared to the difference in FM-CE estimates.
After testing different $m_{\rm dilute}$ structures, we next consider DC models with atmospheric metallicities set to one and three times the stellar value (DCA and DCA3, respectively). The results for the three representative planets are shown in Fig. 3. The complete dataset, including mean values and standard deviations for all structures, is presented in Appendix A. Since $Z_{\rm atm}$ (and thus, the opacity effect) is the varying parameter, DCA and DCA3 directly probe its effect on the $Z$ estimates. As shown in Fig. 3, increasing the atmospheric metallicity systematically increases the inferred $Z$.
@@
In the three cases shown, both DCA and DCA3 yield higher $Z$ than CE. For some planets (such as HAT-P-12b, TOI-2010b, and TOI-2180b), however, DCA results converge with DC models, suggesting that the opacity effects are minor. Overall, Fig. 3 confirms that the atmospheric opacity significantly affects the retrieved heavy-element abundance, whereas density redistribution alone produces comparatively minor shifts. The CEA (bright-red) curves closely track the DCA estimates (especially for TOI-1811b and TOI-5153b), indicating that both models have similar $Z$. Because they use identical opacities and differ only in interior structure (distinct core in CEA vs. composition gradient in DCA), this agreement reinforces that opacity is the dominant lever over the density distribution.
However, while we find that FM models generally lead to higher inferred bulk metallicities, we note that DCA3 models can sometimes lead to even higher estimates, especially for low-$Z$ planets (for example in the cases of TOI-4127b and TOI-2589b). Furthermore, the results are affected by the choice of $Z_{\rm atm}$ since our method does not allow for solutions where the inferred bulk metallicity is lower than the assumed atmospheric metallicity. This limitation particularly affects low-$Z$ planets, as illustrated by the cutoff in the DCA and DCA3 posterior distribution of TOI-5153b shown in Fig. 3. Knowledge of the planetary atmospheric composition is therefore crucial for narrowing the parameter space of possible solutions.
@@ 3.2 Mass-metallicity Relation
After analysing the overall behaviour of the metallicity estimates across different structural models, we next examined the full dataset of 44 planets for possible population-level trends in the inferred composition. In particular, we investigated the heavy-element mass ($M_{Z}$), and the bulk metallicity ($Z$), and their dependency on the planetary mass ($M$).
@@
Figure 4: Relation between the retrieved heavy-element mass $M_{Z}$ and planetary mass $M$, shown for the four main structures tested (CE, FM, DCA, and DCA3). Shown as dashed lines in a grey gradient are curves indicating the trends for fixed planetary compositions (0%, 50%, 90%, and 99% gas composition, respectively). Bayesian fits are shown using a power-law function $M_{Z}=\beta\,M^{\alpha}$. A summary of the fitted parameters and correlation coefficients is presented in Table 1 and Table 2.
@@
To assess these relations, we performed a Bayesian regression. The parameters of the power law $M_{Z}\,[M_{\oplus}]=\beta\,M\,[M_{J}]^{\alpha}$ were estimated using a Markov chain Monte Carlo (MCMC, Python package emcee (Foreman-Mackey et al., 2013), following Müller and Helled (2023) and Howard et al. (2025b)). Posterior distributions of the slope ($\alpha$) and intercept ($\beta$) were sampled from 5,000 chains, taking into account $1\sigma$ uncertainties in both axes.The outputs for each tested structure is reported in Table 1.
All models reveal a positive slope between $M_{Z}$ and $M$, with similar slopes but different intercepts depending on the assumed structure (see Fig. 4). The FM model systematically predicts the highest $M_{Z}$, while CE predicts the lowest (with $\beta\approx 36$). This scaling is broadly consistent with the relation reported by Thorngren et al. (2016), Howard et al. (2025b), and Chachan et al. (2025). Our results suggest a steeper dependence, with DCA3 yielding $\alpha>0.7\pm 0.07$, only marginally inconsistent with Thorngren’s relation of $M_{Z}\propto\sqrt{M}$ when uncertainties are considered.
@@
For this reason, we inspected the reliability of such curves by testing the correlation between $\alpha$ and $\beta$. The covariance analysis shows a weak but positive coupling ($\rho\sim 0.2$ for all the models), indicating that the slope and intercept are not fully independent, and that small shifts in one parameter affect the other. We then also evaluated the fits with a weighted least-squares (WLS) $\chi^{2}$ test. In all cases, $\chi^{2}\gg 1$, suggesting that the power-law relations do not formally explain the data within the quoted errors; therefore, the reliance on such curves should be treated with caution. Not meeting such reliability limits with the $\chi^{2}$ statistic could be caused by (i) noise arising from the scatter in the data, (ii) a large diversity in planetary compositions, or both. In Section D, we further discuss the WSL goodness-of-fit approach.
@@
Figure 5: Relation between retrieved metallicity $Z$ and planetary mass $M$, shown for the four main structures tested (CE, FM, DCA, and DCA3). Bayesian fits are shown using a power-law function, $Z=\beta\,M^{\alpha}$. For additional quantitative results, we refer to Tables 1 and 2.
@@
Since the WSL test revealed the limited reliability of the retrieved power-law fits, it is essential to verify whether the apparent slopes of these relations genuinely reflect an underlying correlation within the data. We further tested the correlations of the dataset for each structural assumption using the non-parametric Kendall’s $\tau$ rank test. To account for observational uncertainties, instead of computing the correlation on the mean values of the data, a bootstrap resampling approach was employed. Here, 5,000 synthetic datasets were generated by drawing from the reported error distributions, and $\tau$ was calculated for each realisation. The median $\tau$ values and associated $p$-values are listed in Table 2. For the $_{Z}MM$ relation, nearly all runs returned $p$-values well below the adopted significance threshold of $p=0.05$, confirming the robustness of the positive correlation.
Table 1: Power-law fit parameters ($y=\beta\,M^{\alpha}$) derived from Bayesian regressions for the four interior-structure models. [^2] Parameters $\alpha$ and $\beta$ correspond to the best-fit power-law relations $M_{Z}=\beta\,M^{\alpha}$ and $Z=\beta\,M^{\alpha}$. Quoted values are posterior medians from Bayesian fits; typical $1\sigma$ uncertainties are $\sigma_{\alpha}\!\simeq\!0.07$ and $\sigma_{\beta}\!\simeq\!4.5$ for $_{Z}MM$, and $\sigma_{\alpha}\!\simeq\!0.07$, $\sigma_{\beta}\!\simeq\!0.014$ for ZZMM.
Model (a) $_{Z}MM$ (b) ZZMM
$\boldsymbol{\alpha}$ $\boldsymbol{\beta}$ $\boldsymbol{\alpha}$ $\boldsymbol{\beta}$
CE 0.67 $\pm$ 0.11 35.9 $\pm$ 4.5 0.33-0.33 $\pm$ 0.12 0.113 $\pm$ 0.014
DCA 0.70 $\pm$ 0.10 43.1 $\pm$ 4.4 0.30-0.30 $\pm$ 0.10 0.136 $\pm$ 0.015
DCA3 0.72 $\pm$ 0.07 56.0 $\pm$ 4.2 0.28-0.28 $\pm$ 0.07 0.176 $\pm$ 0.014
FM 0.65 $\pm$ 0.07 59.6 $\pm$ 4.4 0.35-0.35 $\pm$ 0.07 0.187 $\pm$ 0.014
@@
Table 2: Kendall’s $\tau$ coefficients and associated $p$-values for four structural models (CE, DCA, DCA3, and FM) across the two tested relations, (a) $_{Z}MM$ and (b) ZZMM.
Model (a) $_{Z}MM$ (b) ZZMM
$\boldsymbol{\tau}$ $p$-value $\boldsymbol{\tau}$ $p$-value
CE 0.514 $1.2\times 10^{-6}$ 0.273-0.273 $1.0\times 10^{-2}$
DCA 0.612 $7.6\times 10^{-9}$ 0.245-0.245 $2.1\times 10^{-2}$
DCA3 0.660 $4.4\times 10^{-10}$ 0.314-0.314 $3.1\times 10^{-3}$
FM 0.660 $4.4\times 10^{-10}$ 0.397-0.397 $1.8\times 10^{-4}$
@@
Next, we examined the relation between $Z$ and $M$ (Fig. 5). In contrast to the $_{Z}MM$ trend, the ZZMM relation shows a clear negative correlation. The ranking of models remains similar: FM predicts the highest average $Z$, followed by DCA3, DCA, and CE. Compared to the $_{Z}MM$ relation, the data are more scattered, particularly at lower planetary masses. Kendall’s $\tau$ again indicates statistically significant negative correlations, although the correlation strength is weaker. A portion of the reduced significance may reflect mathematical coupling from dividing by the xx–axis variable ($Z=M_{Z}/M$), which weakens the apparent correlation and inflates $p$-values. Median $p$-values remain generally below 0.05, but with a modest fraction of realisations exceeding the threshold ($\sim 2\%$ for FM and DCA3, and $\sim 10$$11\%$ for DCA and CE). Nevertheless, the correlations remain statistically compelling across all models.
Figure 6: Top: Metallicity estimates for adiabatic ($R_{\rho}$ = 0) and super-adiabatic ($R_{\rho}$ = 0.1, 0.4) models. This is shown for four different planetary cases. Bottom: Corresponding temperature profiles over the planet’s normalised mass, using the structures from the median metallicity of the DCA model at a fixed age of 4 Gyr.
@@
A similar relationship was tested using the normalisation to the host star ($ZZ_{\star}$ vs $M$). The results show similar trends (both in best-fit and correlation tests) to those obtained for ZZ/$M$. Such an investigation would have been essential if we had artificially imposed an anti-correlation driven by the stellar metallicity (especially since one of the model assumptions is that the stellar composition reflects that the planet’s atmospheric structure). Although recent studies (Buchhave et al., 2014; Müller and Helled, 2025) have found trends between planetary heavy-element mass and host star chemical composition, the dataset used in this study does not show any correlation between $Z_{\star}$ and $M$.
@@ 4 Discussion
Figure 7: Kendall’s $\tau$ correlation p-value for $M_{Z}/M$ and $Z/M$ with $n$-planets removed from the initial dataset. Planets are removed from the lowest-mass (left) and highest-mass (right) planets. The number of planets being iteratively removed can be seen on the x-axis. The four different models are respectively colour-coded. The horizontal line shows the $p=0.05$ acceptance threshold.
@@ 4.1 Non-adiabatic structure
In our models with a DC (DC, DCA, and DCA3), we assumed an adiabatic temperature gradient. However, the presence of a composition gradient implies that the temperature gradient could differ from the adiabatic one. To assess the effect of a non-adiabatic temperature gradient on the inferred bulk metallicity, we parameterised the deviation from adiabaticity (likely caused by double-diffusive convection (Debras et al., 2021; Leconte and Chabrier, 2012)) as in Howard et al. (2025a), using
@@
For a smaller selection of planets, we simulated the planetary evolution using $R_{\rho}=0.1$ and 0.4 and compared it to the adiabatic case ($R_{\rho}=0$). The inferred bulk metallicities $Z$ are shown in Fig. 6. We find that $R_{\rho}=0.1$ and $R_{\rho}=0.4$ lead to increased values of $Z$ by up to 7 % and 35 %, respectively. The super-adiabatic temperature gradients obtained with $R_{\rho}=0.1$ and $R_{\rho}=0.4$ yield hotter interiors compared to the adiabatic case (see bottom panel of Fig. 6) and naturally lead to an increased amount of heavy elements required to match the planets’ radii. We note that efforts to constrain $R_{\rho}$ for the Solar System’s gas giants have been recently presented (Sur et al., 2025), although the strong degeneracy in this parameter resulted in a broad range of admissible values. Previous studies estimated $R_{\rho}$ for Jupiter and Saturn, suggesting $R_{\rho}$¡0.1 at around 0.05 (Mankovich et al., 2016; Mankovich and Fortney, 2020). This led us to test values of 0.1 and an extreme case of 0.4. The actual value is model-dependent, and efforts to constrain it in giant planets still remains ongoing.
A clear limitation of our models is that convective mixing is not included. This process could affect the thermal evolution and the distribution of heavy elements. For example, Knierim and Helled (2025) showed that for a Jupiter-mass planet neglecting mixing could overestimate the radius by 0.08 $R_{J}$. The values of $R_{\rho}$ we chose were arbitrary, and it is necessary to assess whether DCs with such a temperature structure can remain stable against convection. Future studies should further assess the effect of non-adiabaticity on the mass–metallicity relation and the thermal evolution of gas giants.